OpenStudy (anonymous):

A wheel has a constant angular acceleration of 3.1 rad/s^2. During a certain 7.0 s interval, it turns through an angle of 140 rad. Assuming that the wheel started from rest, how long had it been in motion before the start of the 7.0 s interval?

OpenStudy (anonymous):

tried doing this

OpenStudy (anonymous):

1 eq: w^2 = wo ^2 +2a(theta) 2 eq: w-wo=at but idk I just can't get it right

ganeshie8 (ganeshie8):

you're given, \(\Delta \theta = 140\) \(\Delta t =7 \) \(\alpha = 3.1\) try finding \(\omega_i\) using : \[\Delta \theta = \omega_i \Delta t +\dfrac{1}{2}\alpha \Delta t^2 \]

ganeshie8 (ganeshie8):

once you have \(\omega_i\), you can find the time taken to reach that velocity

OpenStudy (anonymous):

hi hmm okay wait

OpenStudy (anonymous):

I got wi = 9.15

OpenStudy (anonymous):

v=d/t ???

OpenStudy (anonymous):

to get time?

OpenStudy (anonymous):

or what

OpenStudy (anonymous):

After getting wi(9.15) what equation will I use?

ganeshie8 (ganeshie8):

\(v=u+\alpha t\)

ganeshie8 (ganeshie8):

you know that the initial velocity is 0

ganeshie8 (ganeshie8):

\(\alpha =3.1\) \(v = 9.15\)

ganeshie8 (ganeshie8):

\[9.15 = 0+3.1t\] solve \(t\)

OpenStudy (anonymous):

2.9516 thankyou